# Vector Dark Matter with Higgs Portal in Type II Seesaw framework

### v1

@@ III.1 Experimental constraints
This ratio is constrained to be $\leq 10.5\%$ by ATLAS collaboration ATLAS:2022yvh . Here for $M_{Z^{\prime}}<62.5$ GeV, the coupling of Higgs to DM will be constrained by this condition.
$\bullet$ Constraints from $h\to\gamma\gamma$ and $h\to Z\gamma$ : In the SM, due to the absence of tree level coupling, $h\to\gamma\gamma$ takes place at one loop level via the exchange of charged fermions and $W^{\pm}$ bosons. This particular decay mode plays a significant role in the discovery of the SM Higgs at the LHC. Being a loop mediated process, it shows high sensitivity to any new particles that couple to both the SM Higgs and the photon. Therefore, in the type II seesaw model, the existence of new heavy $H^{\pm}$ and $H^{\pm\pm}$ scalars has the potential to supplement the contributions to this particular channel. The impact of this additional contribution on the $\Gamma(h\to\gamma\gamma)$ can vary, either enhancing or reducing it, depending on the specific model parameters. The partial decay width with these spin-0, spin-$1/2$, spin-1 contribution is given by Shifman:1979eb ; Djouadi:2005gj ; Spira:1997dg ; Gunion:1989we ; BhupalDev:2013xol
@@
The first two terms on the right-hand side denote the SM fermions and W boson contribution in the loop whereas the last two terms represent the non-standard contribution coming from the singly charged ($H^{\pm}$) and the doubly charged ($H^{\pm\pm}$) scalars. Except for the coupling and mass dependence, the $H^{++}$ term has a relative factor of four with respect to the $H^{\pm}$ term in the amplitude due to its enhanced charge.
Similarly, the partial decay width of $h\to Z\gamma$ is given by Arbabifar:2012bd ; Carena:2012xa
@@
ATLAS $1.04^{+0.10}_{-0.09}$ ATLAS:2022tnm $2.0^{+1.0}_{-0.9}$ ATLAS:2020qcv
CMS $1.12^{+0.09}_{-0.09}$ CMS:2021kom $2.4^{-0.9}_{+0.9}$ CMS:2022ahq
@@
Figure 1: $h\to\gamma\gamma$ signal strength as a function of triplet scale $m_{\Delta}$. In the left panel, signal strength as a function of $m_{\Delta}$ ($=m_{H^{+}+},m_{H^{+}},m_{H_{1}},m_{A}$) for different values of $\sin\alpha_{2}$ mentioned inset. In the right panel, the same plot for $\sin\alpha_{2}=0.01$ with different values of the mass difference between charged scalars is considered. The other parameters are fixed at $\sin\alpha_{1}$ is $0.008$, $\sin\alpha_{3}=0.001$, $v_{s}=500$ GeV, $m_{S}=200$ GeV taken for the plots.
@@
Figure 2: $h\to Z\gamma$ signal strength as a function of triplet scale $m_{\Delta}$. In the left panel, signal strength as a function of $m_{\Delta}$ ($=m_{H^{+}+},m_{H^{+}},m_{H_{1}},m_{A}$) for different values of $\sin\alpha_{2}$ mentioned inset. In the right panel, the same plot for $\sin\alpha_{2}=0.01$ with different values of the mass difference between charged scalars is considered. The value of $\sin\alpha_{1}$ is $0.008$, $\sin\alpha_{3}=0.001$, $v_{s}=500$ GeV, $m_{S}=200$ GeV taken for the plots.
The partial decay width of $h\to\gamma\gamma$ and $h\to Z\gamma$ in the type II seesaw extended model differs from SM due to the presence of a triplet scalar. In SM, the $W$ loop interferes with the top loop destructively. In the presence of triplet scalars, the doubly charged and singly charged scalars interfere constructively with the $W$ contribution, as a result, the decay width enhances. The effect of triplet scalar in $h\to\gamma\gamma$ and $h\to Z\gamma$ signal has been discussed in Ref. Arbabifar:2012bd ; BhupalDev:2013xol . Here as another CP even scalar from the complex scalar singlet mixes with the Higgs, the coupling of Higgs to the charged scalars is modified with $\lambda_{\Delta S}$, $\lambda_{s\phi}$ along with mixing angles. The signal strength of $h\to\gamma\gamma$ at LHC is given by
@@
In Fig. 1, the signal strength for $h\to\gamma\gamma$ is shown as a function of the triplet mass scale $m_{\Delta}$. In the left panel, $\mu_{\gamma\gamma}$ is plotted as a function of $m_{\Delta}$ for different values of $\sin\alpha_{2}$. In the right panel, the signal strength of $h\to\gamma\gamma$ is plotted as a function of $m_{\Delta}$ for $\sin\alpha_{2}=0.01$ when different mass gaps between the charged scalars are considered. Here we see that as the dominant couplings that contribute to the decay widths heavily depend on the charged scalar masses, consideration of the non-degenerate mass of the triplet scalars provides us with an enlarged parameter space which is consistent with the measurement of signal strength by ATLAS and CMS. Similar treatment for $\mu_{Z\gamma}$ is done in Fig. 2. Here we see that while non-degeneracy of the charged scalars is considered, the model parameter space for $m_{\Delta}$ up to $600$ GeV predicts a signal strength which is within $1\sigma$ deviation of the ATLAS measurement of signal strength and in $2\sigma$ limit of the CMS measurement of signal strength. Our model parameter space lies within $2\sigma$ limit of a recent combined analysis of ATLAS and CMS search data which show $\mu_{Z\gamma}=2.2\pm 0.7$ at the center of mass energy $13$ TeV ATLAS:2023yqk .
@@ III.2 Theoretical constraints
$\bullet$ Perturbative unitarity: It is also important to check wheather the $2\to 2$ scattering amplitudes involving the scalar fields satisfy unitarity constraints. Consequently the parameters of the scalar potential are constrained from above as the following PhysRevD.16.1519 ; PhysRevD.77.095009 ; GAKEROYD2000119
@@ III.3 Effect of constraints on model parameter space
The conditions described above in Sec. III put stringent constraints on the model parameter space. The conditions on quartic couplings of the scalar potential, which are expressed as functions of the physical masses and the mixing angles, set a tight correlation between these external parameters. We, therefore, search for an allowed parameter space on the basis of physical masses and mixing angles in this section.
While trying to find out a region of the parameter space where all $\lambda$ conditions are satisfied, we observed that degeneracy in the masses of triplet scalars ($m_{H^{+}+},m_{H^{+}},m_{H_{1}},m_{A_{0}}$) is preferred in order to satisfy these conditions. We first search for the allowed value of parameters in $m_{\Delta}$ vs $\sin\alpha_{1}$ plane. Fig. 3 shows the allowed parameter space from unitarity and vacuum stability in the $m_{\Delta}$ vs $\sin\alpha_{1}$ plane. Here we consider $\sin\alpha_{2}$ and $\sin\alpha_{3}$ to be zero implying that $S$ is completely decoupled from the triplet and the SM doublet. The magenta (cyan) coloured points correspond to the case when $v_{t}=2$ ($v_{t}=1$) GeV. The choice of $v_{t}$ has been made keeping it consistent with the $\rho$ parameter constraint. We see in this plot that for a specific value of $\sin\alpha_{1}$ corresponding to each value of $v_{t}$, all values of $m_{\Delta}$ are allowed where $\sin\alpha_{1}\sim(2v_{t}/v_{d})$ For other values of $\sin\alpha_{1}$, $m_{\Delta}\leq 300$ GeV. This feature also has been depicted in the Ref. Ghosh:2017pxl . As $m_{H_{1}}\leq 300$ GeV is already excluded for the degenerate triplet dominated scalars from III.1 (shown by the shaded region) and we are interested in the whole mass range of the triplet scalar for the analysis of DM phenomenology, we set $\sin\alpha_{1}$ as 0.008 for $v_{t}=1~{}{\rm GeV}$ for the rest of this article.
@@
In Fig. 4, we analyse the parametric dependence of $\sin\alpha_{3}$ and $m_{\Delta}$ in $m_{S}$ vs $\sin\alpha_{2}$ plane coming from unitarity and vacuum stability constraints. In the left panel, for the scanning of parameter space in the aforementioned plane, $m_{\Delta}$ has been fixed at 450 GeV whereas in the right panel, it has been fixed at 750 GeV. The different colours in the figure correspond to different values of $\sin\alpha_{3}$. From the variation of $\sin\alpha_{3}$, it is evident that with increment in the mixing angle of singlet and triplet scalar, the allowed masses of $H_{2}$ ($m_{S}$) becomes close to the mass scale of $H_{1}$ ($m_{\Delta}$). In the left panel, the allowed masses of $m_{S}$ are near $450$ GeV which is the scale of the triplet mass when $\sin\alpha_{3}=0.01$. The same feature is observed in the right panel too.
In Fig. 5, we observe the effect of mixing angle $\sin\alpha_{3}$ in $|m_{S}-m_{\Delta}|$ vs $m_{S}$ plane from the same theoretical conditions. Points with dark cyan colour (dark red) corresponds to $\sin\alpha_{3}=0.001$ ($\sin\alpha_{3}=0.01$). We see that the unitarity and vacuum stability constraints allow a mass difference $|m_{S}-m_{\Delta}|\sim 100$ GeV when $\sin\alpha_{3}=0.01$ where as for $\sin\alpha_{3}=0.001$, the allowed mass difference is $|m_{S}-m_{\Delta}|\sim 1000$ GeV.
@@
Along with the other free parameters, $v_{s}$ contributes into the aforementioned conditions through $\lambda_{S}$, $\lambda_{s\phi}$ and $\lambda_{S\Delta}$ parameters. The above $\lambda$ parameters are sensitive to the change in $m_{S}$ and $\sin\alpha_{2}$ and that leads us to analyse the effect of $v_{s}$ in $m_{S}$ vs $\sin\alpha_{2}$ plane.
@@
In Fig. 6, allowed parameter space from theoretical constraints is shown in $m_{S}$ vs $\sin\alpha_{2}$ plane. Variation of $v_{s}$ is shown by assigning different colour to them. The allowed mass region for the whole range of mixing angle diminishes with decreasing $v_{s}$. This is due to the fact that as $v_{s}$ increases, the $\lambda$ parameters connected to it become smaller, therefore the $\lambda$ constraints become weaker. On the contrary, smaller values of $v_{s}$ constrains the parameter space of $m_{S}$ more strongly. The shaded region shown in Fig. 4 and Fig. 6 comes from the exclusion limit of the mixing angle of a scalar singlet with SM Higgs from the LHC as mentioned in III.1.
@@ IV DM phenomenology
Now we are equipped to discuss the DM phenomenology in this model. As a consequence of dark charge conjugation symmetry, $Z^{\prime}$ becomes stable and plays the role of DM in this framework. For a sizable value of new gauge coupling $g_{x}$, $Z^{\prime}$ can be present in the thermal bath of the early universe via its interaction (mediated by $S$) with the SM fields. However, this is possible only when $S$ itself is in thermal equilibrium with the SM fields. Thermalisation of $S$ is dominantly controlled by $\lambda_{s\phi}$. The annihilation channels are responsible for keeping $Z^{\prime}$ and $S$ in the thermal bath with the SM are shown in Fig. 7. We have chosen the model parameters in such a way that the $\lambda_{s\phi}\gtrsim 10^{-4}$ and $g_{x}$ vary between $(0-1)$. Such a choice of parameters ensures the thermalisation of $Z^{\prime}$ in the early universe.
@@
where $g_{\star\rho}$ is the total relativistic degrees of freedom at temperature $T$ contributing to energy density and $M_{\rm Pl}=2.4\times 10^{18}$ GeV. $\langle\sigma_{Z^{\prime}Z^{\prime}\to XX}v_{rel}\rangle$ corresponds to the thermal average of annihilation cross section of $Z^{\prime}$ to SM fields and $H_{1},A_{0},H_{2},H^{\pm\pm},H^{\pm}$. The relic density $\Omega_{Z^{\prime}}h^{2}$ and DD cross sections used in this analysis are calculated with the package micrOMEGAs (v5.3.41) Belanger:2006is in conjunction with FeynRules Alloul:2013bka . In the left panel of Fig. 8, we show how the relic abundance depends on $M_{Z^{\prime}}$. The color bar represents the impact of $g_{x}$. $m_{\Delta}$ and $m_{S}$ have been fixed at $400~{}{\rm GeV}$ and $600~{}{\rm GeV}$ respectively. $\sigma_{Z^{\prime}Z^{\prime}\to SMSM}$ is greatly enhanced at $M_{Z^{\prime}}=m_{h}/2,m_{H_{1}}/2$ and $m_{H_{2}}/2$, due to s-channel resonaces (see Fig. 7). Consequently, relic density falls sharply at these values of $M_{Z^{\prime}}$, as shown in Fig. 8. One can also note that relic density decreases with increasing $g_{x}$ which is the usual behaviour of WIMPs.
@@
In the right panel of Fig. 8, we show the effect of mixing angles in the variation of relic density as a function of $M_{Z^{\prime}}$. The masses of $m_{S}$ and $m_{\Delta}$ are set at $400$ GeV and $1000$ GeV respectively. The three coloured lines in the figure correspond to three different choices of mixing angles $\sin\alpha_{2}$ and $\sin\alpha_{3}$. With the increment of the mixing angles $\sin\alpha_{2}$ and $\sin\alpha_{3}$, the $S$ component increases in the physical scalars $h$ and $H_{1}$. Consequently, the rate of annihilation of DM via the mediation of these physical scalars increases resulting in lower relic abundance of $Z^{\prime}$. We see that for every line when $M_{Z}^{\prime}$ crosses the mass of $S$, relic density falls due to the opening up of dominant $Z^{\prime}Z^{\prime}\to H_{2}H_{2}$ annihilation channel. For $\sin\alpha_{2/3}=0.001$, the effect of $Z^{\prime}Z^{\prime}\to H^{++}H^{--}/H^{+}H^{-}/H_{1}H_{1}$ channels are not observed, but for increasing mixing angles, the effect of these channels comes into play. The magenta line which corresponds to mixing angles of $0.01$, shows another low relic region just after $M_{Z^{\prime}}>m_{\Delta}$. But the $\sin\alpha_{2/3}=0.1$ line shows another interesting effect from the annihilation channel $Z^{\prime}Z^{\prime}\to H_{1}~{}H_{2}$ when $2M_{Z^{\prime}}$ becomes larger than the sum of masses of $H_{2}$ anf $H_{1}$ ($m_{\Delta}+m_{S}$). Additional channels start to contribute to the annihilation of $Z^{\prime}$ significantly when we increase the mixing angles.

### v2

@@ III.1 Experimental constraints
This ratio is constrained to be $\leq 10.5\%$ by ATLAS collaboration Aad et al. (2022). Here for $M_{Z^{\prime}}<62.5$ GeV, the coupling of Higgs to DM will be constrained by this condition.
$\bullet$ Constraints from $h\to\gamma\gamma$ and $h\to Z\gamma$ : In the SM, due to the absence of tree level coupling, $h\to\gamma\gamma$ takes place at one loop level via the exchange of charged fermions and $W^{\pm}$ bosons. This particular decay mode plays a significant role in the discovery of the SM Higgs at the LHC. Being a loop-mediated process, it shows high sensitivity to any new particles that couple to both the SM Higgs and the photon. Therefore, in the type II seesaw model, the existence of new heavy $H^{\pm}$ and $H^{\pm\pm}$ scalars has the potential to supplement the contributions to this particular channel. The impact of this additional contribution on the $\Gamma(h\to\gamma\gamma)$ can vary, either enhancing or reducing it, depending on the specific model parameters. The partial decay width with these spin-0, spin-$1/2$, spin-1 contribution is given by Shifman et al. (1979); Djouadi (2008); Spira (1998); Gunion et al. (2000); Bhupal Dev et al. (2013)
@@
The first two terms on the right-hand side denote the SM fermions and W boson contribution in the loop whereas the last two terms represent the non-standard contribution coming from the singly charged ($H^{\pm}$) and the doubly charged ($H^{\pm\pm}$) scalars. Except for the coupling and mass dependence, the $H^{\pm\pm}$ term has a relative factor of four with respect to the $H^{\pm}$ term in the amplitude due to its enhanced charge.
Similarly, the partial decay width of $h\to Z\gamma$ is given by Arbabifar et al. (2013); Carena et al. (2012)
@@
ATLAS $1.04^{+0.10}_{-0.09}$ Aad et al. (2023b) $2.0^{+1.0}_{-0.9}$ Aad et al. (2020)
CMS $1.12^{+0.09}_{-0.09}$ Sirunyan et al. (2021) $2.4^{-0.9}_{+0.9}$ Tumasyan et al. (2023)
@@
Figure 1: $h\to\gamma\gamma$ signal strength as a function of triplet scale $m_{\Delta}$. In the left panel, signal strength as a function of $m_{\Delta}$ ($=m_{H^{\pm\pm}},m_{H^{\pm}},m_{H_{1}},m_{A}$) for different values of $\sin\alpha_{2}$ mentioned inset. In the right panel, the same plot for $\sin\alpha_{2}=0.01$ with different values of the mass difference between charged scalars is considered. The other parameters are fixed at $\sin\alpha_{1}$ is $0.008$, $\sin\alpha_{3}=0.001$, $v_{s}=500$ GeV, $m_{S}=200$ GeV taken for the plots.
@@
Figure 2: $h\to Z\gamma$ signal strength as a function of triplet scale $m_{\Delta}$. In the left panel, signal strength as a function of $m_{\Delta}$ ($=m_{H^{\pm\pm}},m_{H^{\pm}},m_{H_{1}},m_{A}$) for different values of $\sin\alpha_{2}$ mentioned inset. In the right panel, the same plot for $\sin\alpha_{2}=0.01$ with different values of the mass difference between charged scalars is considered. The value of $\sin\alpha_{1}$ is $0.008$, $\sin\alpha_{3}=0.001$, $v_{s}=500$ GeV, $m_{S}=200$ GeV taken for the plots.
The partial decay width of $h\to\gamma\gamma$ and $h\to Z\gamma$ in the type II seesaw extended model differs from SM due to the presence of a triplet scalar. In SM, the $W$ loop interferes with the top loop destructively. In the presence of triplet scalars, the doubly charged and singly charged scalars interfere constructively with the $W$ contribution, as a result, the decay width enhances. The effect of triplet scalar in $h\to\gamma\gamma$ and $h\to Z\gamma$ signal has been discussed in Ref. Arbabifar et al. (2013); Bhupal Dev et al. (2013). Here as another CP even scalar from the complex scalar singlet mixes with the Higgs, the coupling of Higgs to the charged scalars is modified with $\lambda_{\Delta S}$, $\lambda_{s\phi}$ along with mixing angles. The signal strength of $h\to\gamma\gamma$ at LHC is given by
@@
The experimental bounds on the signal strengths are mentioned in Table.1. In Fig. 1, the signal strength for $h\to\gamma\gamma$ is shown as a function of the triplet mass scale $m_{\Delta}$. In the left panel, $\mu_{\gamma\gamma}$ is plotted as a function of $m_{\Delta}$ for different values of $\sin\alpha_{2}$. In the right panel, the signal strength of $h\to\gamma\gamma$ is plotted as a function of $m_{\Delta}$ for $\sin\alpha_{2}=0.01$ when different mass gaps between the charged scalars are considered. Here we see that as the dominant couplings that contribute to the decay widths heavily depend on the charged scalar masses, consideration of the non-degenerate mass of the triplet scalars provides us with an enlarged parameter space which is consistent with the measurement of signal strength by ATLAS and CMS. Another important point to note is that consideration of the singlet scalar introduces an important parameter $\sin\alpha_{2}$ in the signal strength expression which can affect the value of signal strength $\mu_{\gamma\gamma}$ as shown in the right panel of Fig. 1. Similar treatment for $\mu_{Z\gamma}$ is done in Fig. 2. Here we see that while non-degeneracy of the charged scalars is considered, the model parameter space for $m_{\Delta}$ up to $600$ GeV predicts a signal strength that is within $1\sigma$ deviation of the ATLAS measurement of signal strength and in $2\sigma$ limit of the CMS measurement of signal strength. Our model parameter space lies within $2\sigma$ limit of a recent combined analysis of ATLAS and CMS search data which show $\mu_{Z\gamma}=2.2\pm 0.7$ at the center of mass energy $13$ TeV Aad et al. (2024).
@@ III.2 Theoretical constraints
$\bullet$ Perturbative unitarity: It is also important to check wheather the $2\to 2$ scattering amplitudes involving the scalar fields satisfy unitarity constraints. Consequently the parameters of the scalar potential are constrained from above as the following Lee et al. (1977); Aoki and Kanemura (2008); G Akeroyd et al. (2000)
@@ III.3 Effect of constraints on model parameter space
The experimental and theoretical conditions described in Sec. III put stringent constraints on the model parameter space. The aforementioned constraints on quartic couplings, in turn, can be translated on the physical parameters like masses and mixing, and in the following, we present such model parameter space.
The conditions on $\lambda$ parameters prefer near mass degeneracy among the triplet dominated scalars. Fig. 3 shows the allowed parameter space from unitarity and vacuum stability in the $m_{\Delta}$ vs $\sin\alpha_{1}$ plane. Here we consider $\sin\alpha_{2}$ and $\sin\alpha_{3}$ to be zero implying that $S$ is completely decoupled from the triplet and the SM doublet. The magenta (cyan) coloured points correspond to the case when $v_{t}=2$ ($v_{t}=1$) GeV. The choice of $v_{t}$ has been made keeping it consistent with the $\rho$ parameter constraint. We see in this plot that for a specific value of $\sin\alpha_{1}$ corresponding to each value of $v_{t}$, all values of $m_{\Delta}$ are allowed where $\sin\alpha_{1}\sim(2v_{t}/v_{d})$. For other values of $\sin\alpha_{1}$, $m_{\Delta}\leq 300$ GeV. This feature also has been observed in the Ref. Ghosh et al. (2018). As $m_{H_{1}}\leq 300$ GeV is already excluded for the degenerate triplet dominated scalars from charged Higgs searches (see Sec. III.1) (shown by the shaded region) and we are interested in the whole mass range of the triplet scalar for the analysis of DM phenomenology, we set $\sin\alpha_{1}$ as 0.008 for $v_{t}=1~{}{\rm GeV}$ for the rest of this article.
@@
In Fig. 4, we analyse the parametric dependence of $\sin\alpha_{3}$ and $m_{\Delta}$ on $m_{S}$ vs $\sin\alpha_{2}$ plane coming from unitarity and vacuum stability constraints. In the left panel (right), $m_{\Delta}$ has been fixed at 450 GeV (750 GeV). The different colours in the figure correspond to different values of $\sin\alpha_{3}$. It is evident that with higher values of mixing angle $\sin\alpha_{3}$, the allowed masses of $H_{2}$ ($m_{S}$) become close to the mass scale of $H_{1}$ ($m_{\Delta}$). As for example, in the left (right) panel, the allowed masses of $m_{S}$ are close to $450$ ($750$) GeV which is the scale of the triplet mass when $\sin\alpha_{3}=0.01$.
In Fig. 5, we observe the effect of mixing angle $\sin\alpha_{3}$ in $|m_{S}-m_{\Delta}|$ vs $m_{S}$ plane under the same theoretical constraints. Points with dark cyan colour (dark red) correspond to $\sin\alpha_{3}=0.001$ ($\sin\alpha_{3}=0.01$). We see that the unitarity and vacuum stability constraints allow a mass difference $|m_{S}-m_{\Delta}|\lesssim 100$ GeV when $\sin\alpha_{3}=0.01$ whereas for $\sin\alpha_{3}=0.001$, the allowed mass difference is $|m_{S}-m_{\Delta}|\lesssim 1000$ GeV.
@@
Now we will study the effect of $v_{s}$ on our model parameter space.
@@
In Fig. 6, allowed parameter space from theoretical constraints is shown in $m_{S}$ vs $\sin\alpha_{2}$ plane. The allowed region in $m_{S}$ diminishes with smaller values of $v_{s}$ for all possible values of mixing angle, $\sin\alpha_{2}$. $\lambda_{S}$, $\lambda_{s\phi}$ and $\lambda_{S\Delta}$ become smaller as $v_{s}$ increases (see Eqn. A.1). Consequently, the unitarity constraints on such quartic parameters are automatically satisfied for higher values of $v_{s}$. This has been reflected in larger allowed regions in $m_{S}$ vs $\sin\alpha_{2}$ plane for larger values of $v_{s}$. The shaded region shown in Fig. 4 and Fig. 6 comes from the exclusion limit of the mixing angle of a scalar singlet with SM Higgs from the LHC as mentioned in Sec. III.1.
@@ IV DM phenomenology
Now we are equipped to discuss the DM phenomenology in this model. As a consequence of dark charge conjugation symmetry, $Z^{\prime}$ becomes stable and plays the role of DM in this framework. For a sizable value of new gauge coupling $g_{x}$, $Z^{\prime}$ can be present in the thermal bath of the early universe via its interaction (mediated by $S$) with the SM fields. However, this is possible only when $S$ itself is in thermal equilibrium with the SM fields. Thermalisation of $S$ is dominantly controlled by $\lambda_{s\phi}$. The annihilation channels that are responsible for keeping $Z^{\prime}$ and $S$ in the thermal bath with the SM are shown in Fig. 7. We have chosen the model parameters in such a way that the $\lambda_{s\phi}\gtrsim 10^{-4}$ and $g_{x}$ vary between $(0-1)$. Such a choice of parameters ensures the thermalisation of $Z^{\prime}$ in the early universe.
@@
where $g_{\star\rho}$ is the total relativistic degrees of freedom at temperature $T$ contributing to energy density and $M_{\rm Pl}=2.4\times 10^{18}$ GeV. $\langle\sigma_{Z^{\prime}Z^{\prime}\to XX}v_{rel}\rangle$ corresponds to the thermal average of annihilation cross section of $Z^{\prime}$ to SM fields and $H_{1},A_{0},H_{2},H^{\pm\pm},H^{\pm}$. The relic density $\Omega_{Z^{\prime}}h^{2}$ and DD cross sections used in this analysis are calculated with the package micrOMEGAs (v5.3.41) Belanger et al. (2007) in conjunction with FeynRules Alloul et al. (2014). In the left panel of Fig. 8, we show how the relic abundance depends on $M_{Z^{\prime}}$. The color bar represents the impact of $g_{x}$. $m_{\Delta}$ and $m_{S}$ have been fixed at $400~{}{\rm GeV}$ and $600~{}{\rm GeV}$ respectively. $\sigma_{Z^{\prime}Z^{\prime}\to SMSM}$ is greatly enhanced at $M_{Z^{\prime}}=m_{h}/2,m_{H_{1}}/2$ and $m_{H_{2}}/2$, due to s-channel resonances (see Fig. 7). Consequently, relic density falls sharply at these values of $M_{Z^{\prime}}$, as shown in Fig. 8. One can also note that relic density decreases with increasing $g_{x}$ due to the enhanced annihilation rate of DM which is the usual behaviour of WIMPs.
@@
In the right panel of Fig. 8, we show the effect of mixing angles in the variation of relic density as a function of $M_{Z^{\prime}}$. The masses of $m_{S}$ and $m_{\Delta}$ are set at $400$ GeV and $1000$ GeV respectively. The three coloured lines in the figure correspond to three different choices of mixing angles $\sin\alpha_{2}$ and $\sin\alpha_{3}$. With the increment of the mixing angles $\sin\alpha_{2}$ and $\sin\alpha_{3}$, the $S$ component increases in the physical scalars $h$ and $H_{1}$. Consequently, the rate of annihilation of DM via the mediation of these physical scalars increases resulting in lower relic abundance of $Z^{\prime}$. We see that for every line when $M_{Z}^{\prime}$ crosses the mass of $S$, relic density falls due to the opening up of the dominant $Z^{\prime}Z^{\prime}\to H_{2}H_{2}$ annihilation channel. For $\sin\alpha_{2/3}=0.001$, the effect of $Z^{\prime}Z^{\prime}\to H^{++}H^{--}/H^{+}H^{-}/H_{1}H_{1}$ channels are not observed, but for increasing mixing angles, the effect of these channels comes into play. The magenta line which corresponds to mixing angles of $0.01$, shows another low relic region just after $M_{Z^{\prime}}>m_{\Delta}$. But the $\sin\alpha_{2/3}=0.1$ line shows another interesting effect from the annihilation channel $Z^{\prime}Z^{\prime}\to H_{1}~{}H_{2}$ when $2M_{Z^{\prime}}$ becomes larger than the sum of masses of $H_{2}$ anf $H_{1}$ ($m_{\Delta}+m_{S}$). Additional channels start to contribute to the annihilation of $Z^{\prime}$ significantly when we increase the mixing angles.
In Fig. 9, we show a comparison between the standard $U(1)_{X}$ extended vector dark matter model and our scenario from a dark matter perspective. The scalar sector of the former model consists of the SM Higgs doublet and a scalar singlet whereas our scenario accommodates another $SU(2)_{L}$ triplet scalar along with the above-mentioned particles. In Fig. 9, in the left panel, allowed parameter space for correct relic abundance is shown for a fixed mass of $m_{S}$ and $m_{\Delta}$ in $M_{Z^{\prime}}$ vs $g_{x}$ plane. The blue points correspond to the standard VDM with only an extra singlet scalar whereas the red points correspond to our scenario (with an additional triplet). The first two resonance funnels in the plot appear due to SM Higgs and singlet scalar resonance which are the same for the two scenarios. However, when $M_{Z^{\prime}}>700$ GeV, due to the opening of the extra channel $Z^{\prime}Z^{\prime}\to S\Delta$ in our scenario, relic density is satisfied for a lesser value of $g_{x}$. So the allowed range of $g_{x}$ for a particular value of $M_{Z^{\prime}}$ increases. That result is also reflected in the right panel. The allowed parameter space from relic density constraint is shown by dark blue colored points (dark red points) for the minimal vector dark matter model (our scenario). We observe that the parameter space is larger for our case allowing lower values of $g_{x}$ for higher DM mass ranges. This is due to the fact that adding an extra triplet Higgs in the scalar sector increases the parameter space by adding an extra resonance funnel region (see Fig.8) along with providing new annihilation channels. $\sin\alpha_{3}=0.01$ and $\sin\alpha_{2}=0.1$ is considered for both the plots.
Refer to caption             Refer to caption
Figure 9: Parameter space from relic density constraint in $M_{Z^{\prime}}$ vs $g_{x}$ plane. In the left panel, parameter space for correct relic abundance is shown for a fixed mass of $m_{S}$ and $m_{\Delta}$ (mentioned inset). The dark blue points correspond to the conventional vector dark matter model which has an extra SM singlet in the scalar sector whereas the blue points correspond to our scenario which is extended by another $SU(2)_{L}$ scalar triplet in comparison to the former one. In the right panel, a similar plot has been shown when $m_{S}\in(200-3000)$ GeV and $m_{\Delta}\in(400-3000)$ GeV. $\sin\alpha_{2}=0.1$ is considered for both panels.